# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

### v1

@@ Abstract
Recent alignment methods for large language models, including PPO, DPO, and IPO, are often presented as distinct algorithms. In this work, we show that many of these approaches implicitly conflate two fundamental and independent design choices: (i) the sampling geometry, which determines which samples dominate the gradient signal, and (ii) the optimization geometry, which determines how deviations in value are penalized.
We formalize this observation by expressing alignment as the minimization of a generalized distance between policy energy and target energy, parameterized by an $\alpha$-divergence-based sampling weight and a Bregman divergence-based value metric. We demonstrate that the commonly used KL divergence induces an exponential penalty on unbounded value signals, leading to numerical instability and vanishing gradients in high-confidence regimes.
To address this issue, we propose Orthogonalized Policy Optimization (OPO), a framework that explicitly decouples sampling geometry from optimization geometry. By combining $\alpha$-weighted importance sampling with a $\chi^{2}$-induced quadratic regularization in ratio coordinates, OPO yields a simple and well-conditioned objective with linear gradient dynamics. This formulation maintains stable optimization while preserving peak-seeking behavior, and avoids gradient saturation even when model confidence is high. Our analysis positions OPO as a unifying perspective on existing alignment methods and provides a principled foundation for robust reasoning-oriented training.
@@ 1 Introduction
The field of Large Language Model (LLM) alignment has seen a proliferation of objective functions—PPO [2], DPO [1], IPO [7], GRPO [4]—each demonstrating effectiveness in specific settings. While these methods appear distinct, we observe that they share a common underlying structure that can be decomposed along two independent axes.
Observation: Implicit Conflation of Two Design Choices.
We identify two fundamental degrees of freedom in alignment objectives:
  1. 1.
    Sampling Geometry: Which samples should dominate the gradient? This determines whether optimization is "average-case" (covering the target distribution) or "peak-seeking" (concentrating on high-reward modes).
  2. 2.
    Optimization Geometry: How should deviations in value be penalized? This determines the functional form of the gradient with respect to the policy’s value signal.
In existing methods, these two choices are often coupled through a single divergence measure (typically KL). This implicit conflation can lead to undesirable interactions: for instance, adjusting the KL coefficient to control exploration simultaneously alters the curvature of the optimization landscape.
Diagnosis: The KL-Induced Exponential Penalty.
The standard KL-regularized RL objective induces an exponential penalty on value signals via the log-sum-exp conjugate. While this is well-suited for probability distributions, the Advantage signal in RLHF is an unbounded real value. We show that this mismatch leads to:
  • Numerical Instability: The exponential term can grow unboundedly.
  • Gradient Saturation: In practice, KL-regularized preference objectives often reduce to sigmoid/logistic forms (e.g., DPO), which exhibit gradient saturation for large positive logit margins.
Contribution: Orthogonalized Policy Optimization (OPO).
We propose Orthogonalized Policy Optimization (OPO), a framework that explicitly decouples the two design axes. The key insight is to work in ratio coordinates rather than log-ratio coordinates, where the Pearson $\chi^{2}$ divergence naturally induces a quadratic penalty. The OPO objective is:
$\mathcal{L}_{\text{OPO}}=-\sum_{y\in S}\omega_{\alpha}(y)\cdot v_{\theta}(y)+\frac{\mu}{2}\sum_{y}v_{\theta}(y)^{2},\quad v_{\theta}(y)=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)}-1$ (1)
where $\omega_{\alpha}(y)$ is an $\alpha$-divergence-derived importance weight controlling sampling geometry, and $\frac{\mu}{2}v^{2}$ is the Pearson $\chi^{2}$-induced quadratic regularization controlling optimization geometry. Intuitively, $\omega_{\alpha}$ increases the influence of rare high-quality samples while retaining a controllable interpolation to uniform weighting.
This decoupling allows independent control over "what to learn" (via $\alpha$) and "how to penalize" (via $\mu$), avoiding the entangled dynamics of KL-based methods. We show that this formulation yields a well-conditioned objective with a unique stable equilibrium and linear gradient dynamics.
@@ 3 Theoretical Framework: Two Independent Design Axes
3.2 The Generalized Alignment Objective
We express the general alignment objective as:
$\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\text{sample}}}\left[w_{\alpha}(y)\cdot\mathcal{D}_{\phi}\Big(\text{deviation}_{\theta}(y),\text{target}(y)\Big)\right]$ (4)
This formulation separates two independent design choices:
@@ 3.2 The Generalized Alignment Objective
    Axis 1: Sampling Geometry ($\alpha$): The weight $w_{\alpha}(y)$ determines how samples are weighted in the gradient.
@@
    Axis 2: Optimization Geometry ($\phi$): The distance $\mathcal{D}_{\phi}$ and the choice of deviation coordinates determine the penalty structure.
3.3 Axis 1: Sampling Geometry via $\alpha$-Divergence
The weight $w_{\alpha}(y)$ controls the effective distribution from which the model learns. The parameter $\alpha\in[0,1]$ interpolates between two regimes:
  • $\alpha=1$ (Mode-Covering): $w_{\alpha}(y)\approx 1$. All samples contribute equally. This corresponds to supervised fine-tuning (SFT) behavior.
  • $\alpha\to 0$ (Peak-Seeking): This corresponds to peak-seeking behavior, often yielding exponentially amplified weights under standard RL formulations. High-reward samples dominate the gradient.
Most existing methods implicitly fix $\alpha$. By treating it as an explicit parameter, we enable dynamic interpolation between coverage and exploitation.
3.4 Axis 2: Optimization Geometry via Divergence Choice
The choice of divergence determines the penalty structure. We compare two choices:
KL-Regularized Objectives (Exponential-Type Penalties).
Under standard KL-regularized formulations, the dual induces exponential-type penalties in the log-ratio coordinate via the log-sum-exp operation:
  • Penalty term: $e^{\Delta}$.
  • Gradient scaling: $\nabla\propto e^{\Delta}$.
  • Issue: The gradient grows exponentially with $\Delta$. In practice, KL-regularized preference objectives often reduce to sigmoid/logistic forms (e.g., DPO), which exhibit gradient saturation for large positive logit margins.
$\chi^{2}$ Geometry (Quadratic Penalty in Ratio Coordinates).
The Pearson $\chi^{2}$ divergence naturally induces a quadratic penalty in ratio coordinates:
  • Penalty term: $\frac{1}{2}v^{2}$.
  • Gradient scaling: $\nabla\propto v$.
  • Advantage: The gradient is linear in $v$, providing stable dynamics without saturation.
4 Orthogonalized Policy Optimization (OPO)
Combining the two axes with the $\chi^{2}$ geometry in ratio coordinates, we obtain the OPO objective.
4.1 Pearson $\chi^{2}$ Divergence: A Rigorous Foundation
We first establish the exact relationship between $\chi^{2}$ divergence and quadratic penalties.
Proposition 4.1 (Pearson $\chi^{2}$ induces a quadratic penalty in ratio coordinates).
Let $t(y)=\pi(y)/\pi_{\mathrm{ref}}(y)$ and v(y)=t(y)1v(y)=t(y)-1. The Pearson $\chi^{2}$ divergence satisfies:
$D_{\chi^{2}}(\pi\|\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[(t(y)-1)^{2}\big]=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[v(y)^{2}\big]$ (5)
Therefore, a $\chi^{2}$ trust-region regularization yields a quadratic regularizer in the ratio coordinate $v$.
This is a definitional identity—the Pearson $\chi^{2}$ divergence is the expected squared deviation of the density ratio from unity. This makes $v$ the natural coordinate for $\chi^{2}$-regularized optimization.
4.2 OPO in Ratio Coordinates (Exact Form)
We define Orthogonalized Policy Optimization (OPO) by decoupling: (i) sampling geometry via $\alpha$-weighted importance factors $\omega_{\alpha}(y)$, and (ii) optimization geometry via a Pearson $\chi^{2}$-motivated quadratic penalty in $v_{\theta}$:
@@ 4.2 OPO in Ratio Coordinates (Exact Form)
$\boxed{\mathcal{L}_{\text{OPO}}^{\text{ratio}}=-\sum_{y\in S}\omega_{\alpha}(y)\,v_{\theta}(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[v_{\theta}(y)^{2}\big]}$ (6)
where $v_{\theta}(y)=\pi_{\theta}(y)/\pi_{\mathrm{ref}}(y)-1$ is the centered ratio, and $\omega_{\alpha}(y)$ is an $\alpha$-weighted importance factor. We adopt a generic $\alpha$-divergence-inspired weighting scheme; one practical instantiation is $\omega_{\alpha}(y)=Q(y)\cdot[Q(y)/\tilde{p}_{\theta}(y)]^{1-\alpha}$, though other equivalent constructions are possible.
4.3 Log-Ratio Approximation for LLM Implementation
In practice, LLMs are parameterized in log-probability space, making it natural to work with the log-ratio $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$. We establish the connection via Taylor expansion.
Corollary 4.2 (Local quadratic form in log-ratio space).
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$. Since $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$, for $|\Delta_{\theta}(y)|$ sufficiently small (trust-region regime):
$v_{\theta}(y)=\Delta_{\theta}(y)+O(\Delta_{\theta}(y)^{2}),\qquad v_{\theta}(y)^{2}=\Delta_{\theta}(y)^{2}+O(\Delta_{\theta}(y)^{3})$ (7)
Consequently, Equation˜6 admits the practical approximation:
@@ 4.3 Log-Ratio Approximation for LLM Implementation
which yields linear gradient dynamics in the log-ratio coordinate.
This approximation is valid when the policy remains close to the reference (the trust-region assumption underlying most RLHF methods). In practice, on-policy training with frequent reference updates ensures $|\Delta_{\theta}|$ remains small.
@@ 4.4 Gradient Dynamics
For the ratio-coordinate objective Equation˜6, the gradient w.r.t. $v_{\theta}(y)$ is:
@@ 5 Theoretical Analysis
We now formalize the concept of orthogonality and analyze gradient behavior in high-confidence regimes.
5.1 Orthogonality of Design Axes
Definition 5.1 (Orthogonality of Design Axes).
An alignment objective is orthogonal if the sampling geometry (parameterized by $\alpha$) and the optimization geometry (parameterized by the divergence choice or $\mu$) are decoupled: adjusting one does not implicitly alter the gradient structure determined by the other.
Non-Orthogonality in KL-Based Methods.
In DPO and standard KL-regularized RLHF, the optimal policy has the form $\pi^{*}\propto\pi_{\mathrm{ref}}\exp(r/\beta)$. This introduces $\exp$ into both:
  • The effective sample weight (via the importance ratio).
  • The gradient penalty (via the softmax derivative).
Consequently, changing $\beta$ (intended to control optimization aggressiveness) simultaneously alters the effective sample weighting. The two axes are entangled.
Orthogonality in OPO.
In OPO (Equation˜6), the gradient is $\nabla=-\omega_{\alpha}+\mu v$. We observe:
  • Changing $\alpha$ modifies $\omega_{\alpha}$ but leaves the penalty structure ($\mu v$) unchanged.
  • Changing $\mu$ modifies the regularization stiffness but leaves the importance weights ($\omega_{\alpha}$) unchanged.
This separation allows practitioners to independently tune "which samples matter" and "how strongly to regularize."
5.2 Gradient Behavior in High-Confidence Regimes
A key motivation for OPO is addressing the gradient vanishing issue in reasoning tasks.
The Saturation Problem in Sigmoid-Based Losses.
In DPO and similar methods, the gradient is proportional to:
$\nabla_{\text{DPO}}\propto\sigma(\Delta)(1-\sigma(\Delta))$ (11)
where $\sigma$ is the sigmoid function. As $\Delta\to\infty$ (high confidence), $\sigma(\Delta)\to 1$ and $\nabla\to 0$. This can cause optimization to stall even when further improvement is possible.
Non-Saturating Linear Drive in OPO.
In OPO, the gradient is:
$\nabla_{\text{OPO}}=-\omega_{\alpha}+\mu v$ (12)
Even as $v$ grows large, the gradient remains non-zero as long as $v<\omega_{\alpha}/\mu$. This provides a non-saturating linear drive toward the target, avoiding the saturation that plagues KL-based methods.
Remark 5.2 (Intuition).
Informally, KL-based methods act like a "wall" that becomes harder and harder to push as confidence increases. OPO acts like a "spring" whose restoring force is proportional to displacement, providing consistent feedback throughout training.

### v2

@@ Abstract
Large Language Model (LLM) alignment objectives are often presented as a collection of distinct algorithms—PPO, DPO, IPO, and their variants—each motivated by different derivations. In this work, we argue that this diversity obscures a simpler underlying structure. At a fundamental level, alignment objectives involve two independent design choices: (i) how training signals are sampled and weighted, and (ii) how deviations from a reference policy are geometrically penalized. Existing methods typically entangle these choices through a single divergence, most commonly the Kullback–Leibler (KL) divergence.
We show that this entanglement is not merely a modeling convenience but a source of systematic instability. When the same divergence simultaneously determines sample weighting and optimization curvature, adjusting one aspect (e.g., exploration strength) inevitably alters the other (e.g., gradient geometry). This coupling is particularly problematic in preference-based reinforcement learning, where advantage signals are unbounded and high-confidence regimes are common.
We propose a simple but structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters only as a linear driving force, while optimization geometry is determined independently by a mirror map. This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space. The resulting objective admits a closed-form solution, linear (non-saturating) gradient dynamics, and a well-conditioned trust region, while remaining fully compatible with standard LLM training pipelines.
@@ 1 Introduction
Aligning Large Language Models (LLMs) with human preferences has become a cornerstone of modern AI development. The standard paradigm, Reinforcement Learning from Human Feedback (RLHF) [5, 6], typically employs algorithms like Proximal Policy Optimization (PPO) [2] or Direct Preference Optimization (DPO) [1] to optimize a policy against a reward model or preference dataset. despite their formulation differences, these methods share a common structural foundation: they all rely on Kullback–Leibler (KL) divergence to constrain the policy update, preventing it from deviating excessively from a reference distribution.
While empirically effective, KL-regularized objectives exhibit distinct limitations in reasoning-intensive domains where high confidence is required. As the policy improves and assigns high probability to correct reasoning chains, the KL penalty—which induces an exponential geometry in the log-probability space—often dominates the learning signal. This manifests as gradient saturation, where the driving force for further improvement vanishes exponentially as the model becomes confident. Consequently, training often plateaus prematurely, behavior commonly attributed to "over-regularization" but which we argue is intrinsic to the chosen geometry.
We contend that this issue stems from a structural conflation in existing objective designs. Fundamentally, alignment involves two independent design choices: sampling geometry, which determines the effective weighting of training examples (e.g., whether to focus on high-advantage samples), and optimization geometry, which determines the curvature of the update step (e.g., how to measure distance in the policy space). In standard KL-based methods, a single divergence term dictates both, coupling the exploration strength with the optimization stability. Adjusting one inevitably perturbs the other, creating a dilemma where aggressive sampling destabilizes training, while stable optimization stifles exploration.
To resolve this, we propose a unifying framework based on Orthogonal Mirror Descent. By viewing alignment as a mirror descent process, we can explicitly decouple these two axes: the sampling geometry acts as a linear driving force defined by an advantage-weighted distribution, while the optimization geometry is independently governed by a Bregman divergence (mirror map). This perspective reveals that the saturation observed in DPO and PPO is not a necessary feature of alignment, but a specific consequence of using the Negative Entropy mirror map (which induces KL divergence).
Guided by this framework, we introduce Orthogonalized Policy Optimization (OPO), a new alignment objective derived by selecting the Euclidean mirror map in the space of likelihood ratios. Unlike the exponential geometry of KL, the Euclidean geometry induces a Pearson $\chi^{2}$ trust region, leading to a quadratic regularizer that maintains a linear, non-saturating driving force even in high-confidence regimes. This formulation cleanly separates the sampling intensity (controlled by an $\alpha$-divergence parameter) from the regularization strength (controlled by a stiffness parameter), allowing for robust, continuous improvement.
Our contributions are as follows:
  • We identify the implicit coupling of sampling and optimization geometries in KL-based alignment methods as a root cause of gradient saturation.
  • We propose the Orthogonal Mirror Descent framework, which theoretically decouples these design axes.
  • We derive Orthogonalized Policy Optimization (OPO), a simple, closed-form objective based on Euclidean geometry that ensures linear gradient dynamics and global contraction.
  • We empirically demonstrate that OPO outperforms strong baselines like DPO and GRPO on mathematical reasoning tasks, validating the benefits of non-saturating gradients.
@@ 3 Theoretical Framework: Two Independent Design Axes
4 Theoretical Framework: Orthogonal Mirror Descent
We present a unified framework that fundamentally decouples the preference sampling geometry (what to learn) from the optimization geometry (how to learn). We formulate policy alignment as a functional Orthogonal Mirror Descent problem.
4.1 The General Framework
We view alignment as an iterative projection process. At each step $k$, we seek a policy $\pi$ that maximizes the expected return under a specific sampling distribution $\rho_{\alpha}$, subject to a trust-region constraint defined by a divergence $D_{\Psi}$:
$\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{y\sim\rho_{\alpha}}\left[A(y)\log\pi(y)\right]-\frac{1}{\eta}D_{\Psi}(\pi\|\pi_{k})$ (3)
This formulation cleanly orthogonizes two design axes:
@@ 4.1 The General Framework
    Sampling Geometry ($\rho_{\alpha}$): Determines the effective weighting of samples (the linear driving force).
@@
    Optimization Geometry ($D_{\Psi}$): Determines the mirror map (regularization/curvature) used to project the update.
4.2 Axis 1: Sampling Geometry Specification
The sampling distribution $\rho_{\alpha}$ controls the exploit-explore tradeoff in the data. We define the preference weight $\omega_{\alpha}(y)$ generally as an $\alpha$-parameterized linear signal that reweights the advantage $A(y)$.
$\omega_{\alpha}(y)\propto\text{Projection}_{\alpha}(A(y),\pi_{\mathrm{ref}}(y))$ (4)
Specific constructions include the $\alpha$-divergence optimal target or other heavy-tailed reweighting schemes. This explicit definition avoids ambiguity: $\omega_{\alpha}$ represents the abstract "force" pulling the policy towards high-reward regions, decoupled from the optimization dynamics.
4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
The core innovation of OPO is the choice of the mirror map. Instead of the standard Negative Entropy map (which leads to KL divergence and exponential updates), we choose the Euclidean Mirror Map in the space of likelihood ratios $v=\pi/\pi_{\mathrm{ref}}-1$:
$\Psi(v)=\frac{1}{2}\|v\|_{L^{2}(\pi_{\mathrm{ref}})}^{2}=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}[v(y)^{2}]$ (5)
The induced Bregman divergence is the Pearson $\chi^{2}$ divergence:
$D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\left[\left(\frac{\pi}{\pi_{\mathrm{ref}}}-\frac{\pi_{k}}{\pi_{\mathrm{ref}}}\right)^{2}\right]=D_{\chi^{2}}(\pi\|\pi_{k})$ (6)
This choice transforms the alignment problem into a quadratic optimization in ratio space.
4.4 Orthogonalized Policy Optimization (OPO)
Substituting the Euclidean geometry into our framework yields the OPO objective:
@@ 4.4 Orthogonalized Policy Optimization (OPO)
$\boxed{\mathcal{L}_{\text{OPO}}=-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\omega_{\alpha}(y)\,v(y)\right]+\frac{\mu}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[v(y)^{2}\right]}$ (7)
where $\mu=1/\eta$ is the stiffness parameter. The solution is available in closed form: $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
Remark 4.1 (Analogy with Quantum Amplitudes).
A physical analogy elucidates the advantages of the Euclidean geometry.
  • Ratio Space as Amplitude: In quantum mechanics, superpositions are linear in amplitude ($\psi$). Similarly, OPO operates in the ratio space $v$, allowing preference signals to superpose linearly without saturation.
  • KL Space as Probability: KL-based methods operate in log-probability space. Analogous to adding energies in a Boltzmann distribution, this enforces exponential scaling, which leads to "probability collapse" or saturation when signals are strong.
By working in "amplitude" (ratio) space with quadratic "energy" (regularization), OPO maintains a non-saturating linear response field, crucial for long-chain reasoning where we must avoid premature collapse to a single path.
4.5 Log-Ratio Approximation
For implementation in LLMs (which output logits), we use the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$. As shown in Appendix A, for small trust regions, $v\approx\Delta$.
@@ 4.5 Log-Ratio Approximation
This retains the linear dynamics of the Euclidean geometry while remaining compatible with standard gradient-based training.
Remark 4.2 (On-Policy Anchoring).
We set $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (on-policy anchoring), ensuring the trust region condition $\Delta\approx 0$ holds locally, justifying the approximation.
@@ 4.6 Gradient Dynamics
For the ratio-coordinate objective Equation˜7, the gradient w.r.t. $v_{\theta}(y)$ is:
@@ 5 Theoretical Analysis
We formlize the decoupling of design axes and analyze the distinct regularization properties of OPO.
5.1 Decoupled Sampling and Optimization Geometry
We analyze the optimization dynamics in the function space of ratio outputs $v$. This perspective captures the intrinsic objective geometry. We first formalize the connection to parameter space dynamics.
Lemma 5.1 (Parameter Space Gradient Projection).
Let $\mathcal{L}(v_{\theta})$ be a functional of the policy output $v_{\theta}(y)$. The gradient in parameter space $\theta$ is the projection of the function-space gradient $\nabla_{v}\mathcal{L}$ onto the tangent space of the policy manifold:
$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\nabla_{v}\mathcal{L}(v_{\theta}(y))\cdot\nabla_{\theta}v_{\theta}(y)\right]$ (11)
Thus, the geometric properties of $\nabla_{v}\mathcal{L}$ directly determine the driving force in parameter space, modulated by the policy’s local sensitivity $\nabla_{\theta}v_{\theta}$.
With this link established, we characterize the geometry of $\mathcal{L}(v)$.
Theorem 5.2 (Decoupled Sampling and Optimization Geometry).
Consider the OPO objective $\mathcal{L}(v)=-\sum_{y}\omega_{\alpha}(y)v(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]$. In the space of ratio functions $v\in L^{2}(\pi_{\mathrm{ref}})$:
  1. 1.
    The first-order driving force depends on the sampling parameter $\alpha$:
    $\nabla_{v}\mathcal{L}=-\omega_{\alpha}(y)+\mu v(y)$ (12)
  2. 2.
    The second-order curvature is independent of $\alpha$:
    $\nabla^{2}_{v}\mathcal{L}=\mu I$ (13)
Proof.
Direct differentiation yields a constant Hessian $\mu I$, which is independent of $\alpha$. ∎
This structural decoupling guarantees stable, predictable convergence dynamics in the function space.
Corollary 5.3 (Global Contraction and Linear Convergence).
Gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
$v_{k+1}-v^{*}=(1-\eta\mu)(v_{k}-v^{*})$ (14)
where $v^{*}=\omega_{\alpha}/\mu$. For step size $0<\eta<2/\mu$, this system exhibits global contraction to the unique equilibrium. The contraction rate $|1-\eta\mu|$ depends solely on the optimization stiffness $\mu$, confirming that convergence speed is structurally decoupled from the target distribution sharpness $\alpha$.
We use the term “orthogonalized” to emphasize this control decoupling.
Remark 5.4 (Contrast with KL-Based Methods).
In KL-regularized objectives (e.g., DPO), the effective loss locally resembles a logistic function $\mathcal{L}\approx-\log\sigma(\text{margin})$. The local curvature scales as $\beta^{2}\sigma(m)(1-\sigma(m))$, where $m$ is the logit margin. This curvature depends on both the temperature $\beta$ and the current data margin $m$. Thus, changing $\beta$ simultaneously alters both the gradient magnitude and the local stability profile, entangling the two design axes. OPO’s constant curvature $\mu$ avoids this data-dependent instability.
5.2 Comparison with Parametric L2 Regularization
A natural question is whether OPO differs fundamentally from standard Policy Gradient with L2 parameter regularization (weight decay).
$\mathcal{L}_{\text{L2-PG}}=-\mathbb{E}[A\log\pi_{\theta}]+\frac{\lambda}{2}\|\theta-\theta_{0}\|^{2}$ (15)
While both methods induce a quadratic penalty, they operate in fundamentally different spaces:
  1. 1.
    Functional vs. Parametric Prior: OPO’s penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ acts in the function space of probability ratios. It penalizes deviations only where the policy assigns probability mass. L2 parameter regularization acts in the weight space. Two different policies can have the same L2 distance, and small weight changes can cause large output deviations (and vice-versa).
  2. 2.
    Adaptive Constraint: OPO’s regularization provides a “linear restoring force” directly on the probability ratio. If the model is confident (large $v$), the $\chi^{2}$ penalty generates a strong, specific gradient signal to pull that specific output back. Parametric L2 is blind to the output; it uniformly decays all weights regardless of their contribution to the current output confidence.
  3. 3.
    Learning Dynamics: Empirically, Functional $\chi^{2}$ regularization allows the model to shift mass aggressively where advantages are high, while maintaining stability. Weight decay suppresses all weight magnitudes, which can inadvertently hinder the learning of sharp, high-confidence reasoning chains required for math tasks.
5.3 Quantitative Gradient Analysis
We provide a quantitative comparison of the gradient signal in high-confidence regimes, proving the non-saturation property.
Gradient Saturation in DPO.
For logistic-based losses (DPO), the gradient magnitude with respect to the margin $m$ is $|\nabla_{m}\mathcal{L}|=|\beta\sigma(m)(1-\sigma(m))|$. Since $\sigma(1-\sigma)\leq 1/4$:
$|\nabla_{\text{DPO}}|\leq\frac{\beta}{4}$ (16)
Crucially, as the margin $m\to\infty$, $|\nabla_{\text{DPO}}|\to 0$ exponentially.
Linear Gradient Dynamics in OPO.
For OPO, the gradient is $\nabla_{v}\mathcal{L}=\mu v-\omega_{\alpha}$. Let $v^{*}=\omega_{\alpha}/\mu$. The gradient magnitude is exactly proportional to the distance from equilibrium: $|\nabla_{v}\mathcal{L}|=\mu|v-v^{*}|$. Therefore, for any non-equilibrium state ($|v-v^{*}|\geq\delta$):
$|\nabla_{\text{OPO}}|\geq\mu\delta$ (17)
This guarantees a linear driving force that does not vanish arbitrarily, preventing the saturation observed in KL-based methods.
5.4 Algorithm and Flowchart
We provide a visual overview of the OPO process in Figure 1 and the detailed procedure in Algorithm 2.
Current Policy $\pi_{\theta}$ Sampling $(x,y)$ Evaluation $A(y)$ OPO Decomposition 1. Ratio $r=\pi/\pi_{old}$ 2. Reweight $\omega_{\alpha}\leftarrow r^{\alpha}A$ 3. Orth. Penalty $\mu(r-1)^{2}$ Update $\theta$ Next Iteration
Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework. The core innovation lies in the “OPO Decomposition” block (green), which separates the advantage signal (reweighted by $\alpha$) from the trust-region constraint (quadratic penalty $\chi^{2}$), ensuring linear non-saturating gradients.
Algorithm 1: Orthogonalized Policy Optimization (OPO)  Input: Initial policy $\pi_{\theta}$, hyperparameters $\alpha,\mu,\eta$ for each iteration do 1. Set Reference: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ (on-policy anchoring) 2. Collect Rollouts: Sample $(x,y)\sim\pi_{\theta}$, compute rewards $R(x,y)$ 3. Compute Advantages: $A(y)\leftarrow R(y)-\bar{R}$ (group-normalized) 4. Compute Weights: $\omega_{\alpha}(y)\leftarrow A(y)\cdot[\text{reweight by }\alpha]$ (sampling geometry) 5. Compute Ratio: $v_{\theta}(y)\leftarrow\pi_{\theta}(y)/\pi_{\mathrm{ref}}(y)-1$ (or use log-approx) 6. Compute Loss: $\mathcal{L}\leftarrow-\sum_{y}\omega_{\alpha}(y)v_{\theta}(y)+\frac{\mu}{2}\sum_{y}v_{\theta}(y)^{2}$ (optimization geometry) 7. Update Policy: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$ end for
Figure 2: Pseudocode for Orthogonalized Policy Optimization. The algorithm explicitly decouples sampling geometry (step 4) from optimization geometry (step 6).